OpenStudy (anonymous):

Evaluate limit of x(sqrt(1+x^-2)) as x approaches 0+.

OpenStudy (anonymous):

?

OpenStudy (l):

Take all the terms into the radical and simplify. *

OpenStudy (anonymous):

even the x on the outside?

OpenStudy (anonymous):

The answer I get is 1, but the answer is -1

OpenStudy (l):

x . sqrt(1+1/x^2) = sqrt( x^2 + 1) lim_0 sqrt( x^2 + 1) = 1 It is 1. *

OpenStudy (anonymous):

Maybe the answer key is wrong?

OpenStudy (slaaibak):

xsqrt(1/x^2 +1) = xsqrt(1/x^2(x^2 + 1)) (take out 1/x^-2 as a common factor x*1/x*sqrt(x^2 +1) sqrt(0+1)=1

OpenStudy (slaaibak):

Sorry, I meant 1/x^2 (aka x^-2)

OpenStudy (anonymous):

yeah, I understand. Thank you